OpenStudy (anonymous):

Help! very appreciated. On a certain map, 2 inch represents 16 miles. If Columbus and Deer Creek State Park are 3 1/2" apart on this map, what is their actual distance apart?

OpenStudy (anonymous):

Make a proportion 2/16=3.5/x Cross multiply to get: 2x=(16)(3.5)

OpenStudy (anonymous):

Would my answer be 28?

OpenStudy (anonymous):

Yes! Great job